The temperature of air is 30°C, and the substance cools from 100°C to 70°C in 15 minutes. Find the temperature after 20 min.  


1
Expert's answer
2021-04-15T07:27:10-0400

By Newton’s Cooling Law:

"\lambda\frac{dT}{dt}=30-T"

"t=\intop\frac{\lambda}{30-T}dT=\lambda ln(T-30)+c"


We have: "T=100\degree C" when "t=0"

"c=\lambda ln70"

"t=\lambda(ln70-ln(T-30))=\lambda ln(\frac{70}{T-30})"


We have: "T=70\degree C" when "t=15\ min"

"\lambda=\frac{15}{ln(70/40)}=26.8"


"T=70e^{-t/\lambda}+30=70e^{-t/26.8}+30"


For "t=20\ min" :


"T=70e^{-20/26.8}+30=63.19\degree C"